OpenStudy (anonymous):

Solve using substitution: -2x − y + 1 = 0 2x = -13 − 3y Select one: a. (1, -9) b. (3, -5) c. (-2, 5) d. (4, -7)

OpenStudy (anonymous):

d. (4, -7)

OpenStudy (anonymous):

Thanks :) up for anotheer?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-2x − y + 1 = 0 2x = -13 − 3y or -2x − y = -1 2x + 3y = - 13 add the first eq to the second 2y = -14 y = -7 subst into the first -2x +7 = -1 substr 7 from both sides -2x = -8 x = 4 so (x,Y) = (4, -7) Hence

OpenStudy (anonymous):

d